OpenStudy (anonymous):

Identify the factors of 9x2 + 30x + 25

OpenStudy (anonymous):

do you have any idea on how to approach it? remember: we already worked one before out :)

OpenStudy (anonymous):

one way: use the quadratic formula

OpenStudy (anonymous):

is it 9x^2 + 30x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

take a look at the one we did previously

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

9x^2 X 30x

OpenStudy (anonymous):

I think I forgot it

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

quadratic formula ^

OpenStudy (anonymous):

ax^2 + bx + c 9x^2 + 30x + 25 where a = 9, b = 30, c = 25

OpenStudy (anonymous):

\[x = \frac{ -(30) \pm \sqrt{(30)^2 - 4(9)(25)} }{ 2(9) }\]

OpenStudy (anonymous):

thank you